# LOB-Bench: Benchmarking Generative AI for Finance – an Application to Limit Order Book Data

### v1

@@ Abstract
While financial data presents one of the most challenging and interesting sequence modelling tasks due to high noise, heavy tails, and strategic interactions, progress in this area has been hindered by the lack of consensus on quantitative evaluation paradigms. To address this, we present LOB-Bench, a benchmark, implemented in python, designed to evaluate the quality and realism of generative message-by-order data for limit order books (LOB) in the LOBSTER format. Our framework measures distributional differences in conditional and unconditional statistics between generated and real LOB data, supporting flexible multivariate statistical evaluation. The benchmark also includes commonly used LOB statistics such as spread, order book volumes, order imbalance, and message inter-arrival times, along with scores from a trained discriminator network. Lastly, LOB-Bench contains “market impact metrics”, i.e. the cross-correlations and price response functions for specific events in the data. We benchmark generative autoregressive state-space models, a (C)GAN, as well as a parametric LOB model, and find that the autoregressive GenAI approach beats traditional model classes. Code and generated data are available at: https://github.com/peernagy/lob_bench.
@@ 1 Introduction
Figure 2: Schematic of the LOB.
@@
To address this, we propose a general framework for evaluating the similarity between the distribution induced by generative LOB models and the ground-truth data. At a high level, our unconditional evaluation consists of three steps. We first introduce a set of aggregator functions, $\bm{\Phi}$, which map from high dimensional time-series LOB data into a set of 1d subspaces. Secondly, we compute histograms to estimate distributions for the ground-truth and generated data in these subspaces and, finally, use a distance metric, e.g. L1, to compare differences in these estimates. Some of the aggregator functions chosen are closely inspired by metrics used in literature (Vyetrenko et al., 2021; Paulin, 2019; Chiarella & Iori, 2002; Cont, 2001). They also directly relate to generative adversarial networks, where the discriminator network is equivalent to a worst-case aggregator function for a given generator.
@@
  • Identification of a common failure mode: divergence metrics, computed as distributional errors as a function of unroll step, highlight a prevalent failure mode that can guide research.
@@
  • Transferability to other domains: theoretical framework which is adaptable to other high-dimensional generative time-series tasks beyond LOB data.
We hope our benchmark will provide a much-needed starting point for evaluating GenAI models in finance and allow more machine learning scientists to develop new sequence models for this important and challenging domain. Our code is available at: https://github.com/peernagy/lob_bench.
@@ 2.1 Limit Order Book (LOB)
Later sections of this paper rely on the reader’s understanding of the mechanisms of electronic markets, so we briefly review them here. Public exchanges such as NASDAQ and NYSE facilitate the buying and selling of assets by accepting and satisfying buy and sell orders from multiple market participants. The exchange maintains an order book data structure for each asset traded. The LOB represents a snapshot of the supply and demand for the asset at a given time. It is an electronic record of all the outstanding buy and sell limit orders organized by price levels. A matching engine, such as the price-time priority, is used to pair incoming buy and sell order interest as mentioned in Bouchaud et al. (2018). Order types are further distinguished between limit orders and market orders. A limit order (Figure 2) specifies a price that should not be exceeded in the case of a buy order (bid), or should not be gone below in the case of a sell order (ask). A limit order queues a resting order in the LOB at the corresponding side of the book. A market order indicates that the trader is willing to accept the best price available immediately.
@@ 2.3 Autoregressive LOB models
In machine learning, autoregressive modelling is a key component of language models like GPT. By learning the probability distribution of the next token given the previous tokens, autoregressive language models can generate coherent text (Radford et al., 2019). Cross-entropy is a loss function commonly used to train classification models in deep learning. It measures the dissimilarity between the predicted class probabilities and the true class labels (Goodfellow et al., 2016). Cross-entropy loss is the negative log likelihood of the true class labels under the predicted distribution. Minimizing the cross-entropy is equivalent to maximizing the likelihood of the data (Murphy, 2012). For binary classification, the cross-entropy loss is:
$L=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right]$
where $y_{i}$ is the true label (0 or 1) and $\hat{y}_{i}$ is the predicted probability for the positive class. Cross-entropy loss heavily penalizes confident misclassifications and incentivizes the model to output calibrated probabilities that match the empirical distribution of the classes. Although it is different from the KL divergence, cross-entropy can be expressed as the sum of the entropy of the true distribution and the KL divergence between the true and predicted distributions (Cover & Thomas, 1999).
@@ 3 Related Literature
The LOB plays a crucial role in modern financial markets. With the FI-2010 dataset, Ntakaris et al. (2018) released the first publicly available high-frequency LOB dataset for benchmarking mid-price prediction models. This dataset orders for five stocks on the Nasdaq Nordic market for ten consecutive days, and is pre-processed. Although useful and effective for preliminary tests and comparisons of LOB algorithms, FI-2010 does not allow a comprehensive evaluation of robustness and generalisation ability (Zhang et al., 2019). A similar benchmark for average price and volume prediction in Chinese stock markets is provided by Huang et al. (2021). As with other currently available benchmarks, this work falls short of evaluating GenAI models with a fully distributional lens. Cao et al. (2022) propose a benchmark dataset, which plays a complementary role to LOB-Bench. With DSLOB, they provide a synthetic LOB dataset, generated by a multi-agent simulation with shocks, which generates labelled in- and out-of-distributions samples. In contrast, LOB-Bench does not require training on a specific dataset, and instead focuses on general-purpose model evaluation and comparison.
@@ 4 Evaluation Framework
To evaluate the magnitude of the “autoregressive trap” the benchmark evaluates error divergence of distributions, conditional on the interval of the forecasting step $t\in\mathbb{N}$, for interval limits $a,b\in\mathbb{N}$: $\mathbb{D}\left[p\left\{\Phi(d_{t\in[a,b)})\right\}\mid\mid\hat{p}\left\{\Phi(d_{t\in[a,b)})\right\}\right]$. This allows quantifying distribution shift during inference.
Our framework uses both the L1 norm and the Wasserstein-1 distance as loss metrics. To estimate the L1 norm we first bin the data. As a robust binning algorithm, we use the Freedman-Diaconis rule (Freedman & Diaconis, 1981), which computes the bin width as $2\frac{IQR}{\sqrt[3]{n}}$, where $n$ is the combined sample size of the real and generated data. The $[0,1]$-scaled L1 norm, also called total variation distance, can then be estimated as:
@@
While the L1 measure has the benefit of being bounded in the interval $[0,1]$, the Wasserstein-1 distance, or earth mover’s distance, as proposed by Rubner et al. (2000), has the advantage of being sensitive to the distance between the scores. To make losses between different scoring functions comparable, we mean-variance normalize the data before calculating the Wasserstein-1 distance.
@@ 4.1 Impact response functions
A primary difficulty with historical LOB data is that counterfactual scenarios are impossible to evaluate due to being static not responding additional injected orders. Generative models are therefore a unique opportunity to generate a response to counterfactual scenarios.
It is therefore crucial that such models be evaluated on their ability to provide a realistic response to different events. As an underlying methodology, the seminal work by Eisler et al. (2012) is used as a basis to compare the impact of different event types. This methodology focuses only on the impact of events, which change the price or quantity of the best bid and ask orders (also known as touch orders), which is one of its limitations.
@@
This allows calculation of the response function ((6)). This is calculated empirically using the time average (.$_{T}$) of the change in the sign-adjusted mid-price $p_{t}=\frac{a_{t}+b_{t}}{2}$ following a given event, for different lag-times $l$. The event lag times are chosen to be distributed uniformly on a logarithmic scale between 1 and 200 ticks. The prices are normalized by tick size to enable a comparison between various stocks.
@@ 4.2 Adversarial Measurement
The concept of adversarial measurement is to develop a pre-trained discriminator capable of effectively distinguishing between true and generated trajectories. This discriminator is a binary classifier, generating a probability estimate of a trajectory being real. We use only the orderbook states as an input. The discriminator is trained on two batches of data, each of dimension $(S\times T\times D)$. In this representation, $S$ denotes the number of sequence samples within the batch, $T$ the length of the orderbook sequences, and $D$ is the dimension of the orderbook state representation. Given the sparsity of changes between most orderbook states, we devised an encoding scheme to optimize the discriminator’s performance.
@@
The discriminator aims to find the “worst-case” function $\Phi^{*}$ that maximally separates the real and generated distributions by choosing $\Phi^{*}$ such that it maximizes the divergence between them, i.e., $\Phi^{*}=\arg\max_{\Phi}D[p(\Phi(d)),\hat{p}(\Phi(d)))]$. This $\Phi^{*}$, which can be interpreted as a dimensionality reduction operation on the sequence of order book states $\mathbf{b}\in\mathcal{B}$ to a scalar $s$, $\Phi^{*}:(\mathcal{M}\times\mathcal{B})\mapsto\mathbb{R}$, can be considered to be an adversarial scoring function. The discriminator tries to identify the most glaring flaws and differences between the real and generated samples and distil these to a single dimension.
@@
An orderbook state comprises the price and quantity from the top $n$ price levels on both the bid and ask sides. In this paper, $n=10$ resulting in a dimension $D=40$. Changes in the orderbook state are typically triggered by events that affect a single price-quantity pair. To achieve a more concise, yet informative, representation of the discriminator network, we chose to represent the orderbook based on these changes. Thus the book states $\mathbf{b}\in\mathcal{B}$ and message events $\mathbf{m}\in\mathcal{M}$ map to three-dimensional vectors through $i\in\mathbb{N}$ functions $\Psi_{i}:(\mathcal{M}\times\mathcal{B})\mapsto\mathbb{R}^{3}$. These changes encompass each change in the mid-price, the relative price level where the change occurs, and the corresponding change in quantity. Our discriminator utilizes a 1D convolutional neural network (Conv1D) (LeCun et al., 1995; Kiranyaz et al., 2019) as a feature extractor, followed by an attention mechanism (Vaswani et al., 2017) to capture long-term dependencies across the time steps. Empirical results show that this model, trained and tested on GOOG data from 2023, achieves a Receiver Operating Characteristic (ROC) score of 0.83, indicating that the generated data can be discriminated reasonably accurately. However, the baseline model’s performance for GOOG and INTC was poor, with a discriminator ROC score of around 1, indicating significant room for future model improvement.
@@ 5 LOB-Bench Package
Based on the evaluation framework outlined in section 4, we developed a Python benchmark package, allowing for a convenient and comprehensive evaluation of generated LOB data. The benchmark is highly customizable as scoring functions $\Phi$ can easily be added, removed or modified, and provides a standardized model comparison using the provided default scoring functions. The benchmark reports aggregate model scores by computing the mean, median, and inter-quartile mean (IQM111mean of all values between the 25. and 75. percentile) across all conditional and unconditional scoring functions, along with bootstrapped confidence intervals.
The benchmark performs both unconditional and conditional evaluation of generated data, by computing distributions of statistics of interest conditionally on the value of another statistic. To evaluate the magnitude of the effect of error divergence or “snowballing errors,” distributions are also evaluated conditional on the prediction horizon. Distributional accuracy is measured by computing the L1-norm and Wasserstein-1 distance between the real and generated distributions. Specific supported examples of more complex conditional distributions are the response functions, describing the distribution of events conditional on other events having occurred at a certain prior lag. As these distributions usually have high variance, and to be consistent with the extant literature, we instead measure mean absolute differences in their means for a range of lags to evaluate market impact curves.
@@
Volume per Minute Traded volume in one-second intervals scaled to a minute.
@@
OFI (Up/Stay/Down) OFI (see above) conditional on the next message moving the mid-price Up/Static/Down
The benchmark also evaluates model response functions (6) in aggregate. Individual L1 distances $\Delta R_{\pi}$ are calculated for each lag time and averaged to produce aggregate impact scores.

### v2

@@ Abstract
While financial data presents one of the most challenging and interesting sequence modelling tasks due to high noise, heavy tails, and strategic interactions, progress in this area has been hindered by the lack of consensus on quantitative evaluation paradigms. To address this, we present LOB-Bench, a benchmark, implemented in python, designed to evaluate the quality and realism of generative message-by-order data for limit order books (LOB) in the LOBSTER format. Our framework measures distributional differences in conditional and unconditional statistics between generated and real LOB data, supporting flexible multivariate statistical evaluation. The benchmark also includes commonly used LOB statistics such as spread, order book volumes, order imbalance, and message inter-arrival times, along with scores from a trained discriminator network. Lastly, LOB-Bench contains “market impact metrics”, i.e. the cross-correlations and price response functions for specific events in the data. We benchmark generative autoregressive state-space models, a (C)GAN, as well as a parametric LOB model, and find that the autoregressive GenAI approach beats traditional model classes. Code and generated data are available at: https://lobbench.github.io/.
@@ 1 Introduction
Figure 2: Schematic of the LOB. Not immediately executable orders are placed in First-In, First-Out (FIFO) queues at the specified limit price. Sell limit orders are placed on the Ask side (red) and Buy limit orders on the Bid side (blue).
@@
To address this, we propose a general framework for evaluating the similarity between the distribution induced by generative LOB models and the ground-truth data. At a high level, our unconditional evaluation consists of three steps. We first introduce a set of aggregator functions, $\bm{\Phi}$, which map from high-dimensional time series LOB data into a set of 1d subspaces. Secondly, we compute histograms to estimate distributions for the ground-truth and generated data in these subspaces and, finally, use a distance metric, e.g. $L_{1}$, to compare differences in these estimates. Some of the aggregator functions chosen are closely inspired by metrics used in literature (Vyetrenko et al., 2021; Paulin, 2019; Chiarella & Iori, 2002; Cont, 2001). They also directly relate to generative adversarial networks (Goodfellow et al., 2014), where the discriminator network is equivalent to a worst-case aggregator function for a given generator.
@@
  • Identification of a common failure mode: divergence metrics, computed as distributional errors as a function of unroll step, highlight a prevalent failure mode to guide research.
@@
  • Transferability to other domains: the theoretical framework is adaptable to other high-dimensional generative time series tasks beyond LOB data.
We hope our benchmark will provide a much-needed starting point for evaluating GenAI models in finance and allow more machine learning scientists to develop new sequence models for this important and challenging domain. Our code and additional resources are available at: https://lobbench.github.io/.
@@ 2.1 Limit Order Book (LOB)
Later sections of this paper rely on the reader’s understanding of the mechanisms of electronic markets, so we briefly review them here. Public exchanges such as NASDAQ and NYSE facilitate the buying and selling of assets by accepting and satisfying buy and sell orders from multiple market participants. The exchange maintains an order book data structure for each asset traded. The LOB represents a snapshot of the supply and demand for the asset at a given time. It is an electronic record of all the outstanding buy and sell limit orders organized by price levels. A matching engine, commonly using a price-time priority mechanism, is used to pair incoming buy and sell order interest as mentioned in Bouchaud et al. (2018). Order types are further distinguished between limit orders and market orders. A limit order (Figure 2) specifies a price that should not be exceeded in the case of a buy order (bid), or should not be gone below in the case of a sell order (ask). A limit order queues a resting order in the LOB at the corresponding side of the book. A market order indicates that the trader is willing to accept the best price available immediately.
@@ 2.3 Autoregressive LOB models
In machine learning, autoregressive modelling is a key component of language models like GPT. By learning the probability distribution of the next token given the previous tokens, autoregressive language models can generate coherent text (Radford et al., 2019). Cross-entropy is a loss function commonly used to train classification models in deep learning. It measures the dissimilarity between the predicted class probabilities and the true class labels (Goodfellow et al., 2016). Cross-entropy loss is the negative log likelihood of the true class labels under the predicted distribution. Minimizing the cross-entropy is equivalent to maximizing the likelihood of the data (Murphy, 2012). The cross-entropy loss over a sample of size $N$, with $V$ classes can be expressed as:
$L=-\frac{1}{N}\sum_{i=1}^{N}\sum_{v=1}^{V}y_{i}^{(v)}\log\hat{y}_{i}^{(v)},$
where $y_{i}^{(v)}$ equals 1 if the true class is $v$ and 0 otherwise, and $\hat{y}_{i}^{(v)}$ is the predicted probability for class $v$. Cross-entropy loss heavily penalizes confident misclassifications and incentivizes the model to output calibrated probabilities that match the empirical distribution of the classes. Although it is different from the KL divergence, cross-entropy can be expressed as the sum of the entropy of the true distribution and the KL divergence between the true and predicted distributions (Cover & Thomas, 1999).
@@ 3 Related Literature
The LOB plays a crucial role in modern financial markets. With the FI-2010 dataset, Ntakaris et al. (2018) released the first publicly available high-frequency LOB dataset for benchmarking mid-price prediction models. This pre-processed dataset containst orders for five stocks on the Nasdaq Nordic market for ten consecutive days. Although useful and effective for preliminary tests and comparisons of LOB algorithms, FI-2010 does not allow a comprehensive evaluation of robustness and generalisation ability (Zhang et al., 2019). A similar benchmark for average price and volume prediction in Chinese stock markets is provided by Huang et al. (2021). As with other currently available benchmarks, this work falls short of evaluating GenAI models with a fully distributional lens. Cao et al. (2022) propose a benchmark dataset, which plays a complementary role to LOB-Bench. With DSLOB, they provide a synthetic LOB dataset, generated by a multi-agent simulation with shocks, which generates labelled in- and out-of-distributions samples. In contrast, LOB-Bench does not require training on a specific dataset, and instead focuses on general-purpose model evaluation and comparison.
@@ 4 Evaluation Framework
To evaluate the magnitude of the “autoregressive trap” the benchmark evaluates error divergence of distributions, conditional on the interval of the forecasting step $t\in\mathbb{N}$, for interval limits $a,b\in\mathbb{N}$: $\mathbb{D}\left[p\left\{\Phi(d)\right\}\mid\mid\hat{p}\left\{\Phi(d_{t\in[a,b)})\right\}\right]$. This allows quantifying distribution shift during inference.
Our framework uses both the $L_{1}$ norm and the Wasserstein-1 distance as loss metrics. To estimate the $L_{1}$ norm, we first bin the data. As a robust binning algorithm, we use the Freedman-Diaconis rule (Freedman & Diaconis, 1981), which computes the bin width as $2\frac{IQR}{\sqrt[3]{n}}$, where $n$ is the combined sample size and IQR the inter-quartile range of the real and generated data. The $[0,1]$-scaled $L_{1}$ norm, also called the total variation distance, can then be estimated as:
@@
While the $L_{1}$ measure has the benefit of being bounded in the interval $[0,1]$, the Wasserstein-1 distance, or earth mover’s distance, as proposed by Rubner et al. (2000), has the advantage of being sensitive to the distance between the scores. To make losses between different scoring functions comparable, we mean-variance normalize the data before calculating the Wasserstein-1 distance.
@@
Our methodology formalizes and naturally extends common evaluation practices for synthetic one-dimensional time series, such as financial returns, which typically emphasize distributional similarity. Our framework enables a quantitative assessment of distributional properties in structured high-dimensional time series. By adapting the scoring functions, our approach could also be applied to financial transactions, payment data, streamed price quotes in forex markets, multi-asset limit order books, or decentralized crypto market protocols.
@@ 4.1 Impact Response Functions
A primary difficulty with historical LOB data is that counterfactual scenarios are impossible to evaluate, as the data do not respond to additional injected orders. Generative models are a unique opportunity to generate a response to counterfactual scenarios as they address this limitation.
It is therefore crucial that such models be evaluated on their ability to provide a realistic response to different events. As an underlying methodology, the seminal work by Eisler et al. (2012) is used as a basis to compare the impact of different event types. This methodology focuses only on the impact of events, which change the price or quantity of the best bid and ask orders (also known as touch orders), which also constitutes a key limitation of the method.
@@
This allows calculation of the response function – equation (6). This is calculated empirically using the time average (.$_{T}$) of the change in the sign-adjusted mid-price $p_{t}=\frac{a_{t}+b_{t}}{2}$ following a given event, for different lag-times $l$. The event lag times are chosen to be distributed uniformly on a logarithmic scale between 1 and 200 ticks. The prices are normalized by tick size to enable a comparison between various stocks.
@@ 4.2 Adversarial Measurement
The concept of adversarial measurement involved developing a pre-trained discriminator capable of effectively distinguishing between true and generated trajectories. This discriminator is a binary classifier, generating a probability estimate of a trajectory being real. We only use the orderbook states as input, ignoring the message sequences. The discriminator is trained using two batches of data, each of dimension $(S\times T\times D)$. $S$ denotes the number of sequence samples within the batch, $T$ the length of the orderbook sequences, and $D$ is the dimension of the orderbook state representation.
@@
The discriminator aims to find the “worst-case” function $\Phi^{*}$ that maximally separates the real and generated distributions by choosing $\Phi^{*}$ such that it maximizes the divergence between them, i.e., $\Phi^{*}=\arg\max_{\Phi}D[p(\Phi(d)),\hat{p}(\Phi(d)))]$. This $\Phi^{*}$, which can be interpreted as a dimensionality reduction operation on the sequence of order book states $\mathbf{b}\in\mathcal{B}$ to a scalar $s$, $\Phi^{*}:(\mathcal{M}\times\mathcal{B})\mapsto\mathbb{R}$, can be considered to be an adversarial scoring function. The discriminator attempts to identify the most glaring flaws and differences between the real and generated samples, distilling these into a single dimension.
@@
Given the sparsity of changes between successive orderbook states, we devised an encoding scheme to optimize the discriminator’s performance. An orderbook state comprises the price and quantity from the top $n$ price levels on both the bid and ask sides. In our experiments, $n=10$ resulting in dimension $D=40$. Changes in the orderbook state are typically triggered by events that affect a single price-quantity pair. To achieve a more concise, yet informative, representation of the discriminator network, we chose to represent the orderbook based on these changes. Thus the book states $\mathbf{b}\in\mathcal{B}$ and message events $\mathbf{m}\in\mathcal{M}$ map to three-dimensional vectors through $i\in\mathbb{N}$ functions $\Psi_{i}:(\mathcal{M}\times\mathcal{B})\mapsto\mathbb{R}^{3}$. These changes encompass each change in the mid-price, the relative price level where the change occurs, and the corresponding change in quantity. Our discriminator utilizes a 1D convolutional neural network (Conv1D) (Lecun & Bengio, 1995; Kiranyaz et al., 2019) as a feature extractor, followed by an attention mechanism (Vaswani et al., 2017) to capture long-term dependencies across the time steps. Empirical results show that this model, trained and tested on GOOG data from 2023, achieves a Receiver Operating Characteristic (ROC) score of 0.83, indicating that the generated data can be discriminated reasonably accurately. However, the baseline model’s performance for GOOG and INTC was poor, with a discriminator ROC score of around 1, indicating significant room for future model improvement. High discriminability may result from model errors, as indicated by imperfect model scores (see Section 6 and Figures 13ff). A distributional mismatch in a single scoring function can be sufficient to make fake data identifiable. To mitigate this issue, future research could evaluate adversarial performance by training a discriminator on perturbed data and reporting scores conditioned on the noise level, particularly as models improve on this benchmark.
4.3 Mid-price Prediction
To evaluate the potential value added from additional generated synthetic data, we measure the impact of the additional training data on prediction quality for a simple downstream task. Concretely, we adapt the LOBCAST (Prata et al., 2024) implementation to train an MLP with one million parameters to classify the mid-price movement for different prediction horizons. The MLP is trained to predict three classes: up, down, and stationary. Following Prata et al. (2024), classes are defined based on the movement of the mid-price over a threshold value. For each generative model evaluated, two separate MLPs are trained, one on only real data, and another on real & generated data. F1-scores calculated on held-out real data are then evaluated as a measure of the impact of using the generated data in training.
@@ 5 LOB-Bench Package
Based on the evaluation framework outlined in section 4, we develop a Python benchmark package, allowing for a convenient and comprehensive evaluation of generated LOB data. The benchmark is highly customizable, as scoring functions $\Phi$ can easily be added, removed, or modified, and provides a standardized model comparison using the default scoring functions provided. The benchmark reports aggregate model scores by computing the mean, median, and inter-quartile mean (IQM111mean of all values between the 25. and 75. percentile) across all conditional and unconditional scoring functions, along with bootstrapped confidence intervals.
The benchmark performs both unconditional and conditional evaluation of generated data, by computing distributions of statistics of interest conditionally on the value of another statistic. To evaluate the magnitude of the effect of error divergence or “snowballing errors,” distributions are also evaluated conditional on the prediction horizon. Distributional accuracy is measured by computing the $L_{1}$-norm and Wasserstein-1 distance between the real and generated distributions. Specific supported examples of more complex conditional distributions are the response functions, describing the distribution of events conditional on other events having occurred at a certain prior lag. As these distributions usually have high variance, and to be consistent with the extant literature, we instead measure mean absolute differences in their means for a range of lags to evaluate market impact curves.
@@
Volume per Minute Traded volume in one-second intervals, scaled to a minute.
@@
OFI (Up/Stay/Down) OFI (see above), conditional on the subsequent message’s mid-price move: Up/Static/Down
Figure 4: Model score summaries (lower is better). The LOBS5 model achieves the lowest overall scores. Coletta beats the baseline on the Wasserstein metric, but not for $L_{1}$. Error bars are bootstrapped 99% CIs.
The benchmark also evaluates model response functions (6) in aggregate. Individual $L_{1}$ distances $\Delta R_{\pi}$ are calculated for each lag time and averaged to produce aggregate impact scores.